# FGM Modeling of Thermo-Diffusive Unstable Lean Premixed Hydrogen-Air Flames

### v1

@@ 1 Introduction
Lean laminar TD unstable hydrogen-air flames have been investigated by numerous sources in the literature [1, 2, 3, 4, 5, 6, 7, 8]. Altantzis et al. [1, 2] studied the initial linear growth of superimposed perturbations and the long-term nonlinear evolution of instabilities by 2-D numerical simulations with single-step chemistry and detailed transport for varying domain sizes. They show that positively-curved segments of the flame experience increased reactivity, while the reactivity in negatively-curved segments decreases. Furthermore, that the domain size significantly influences the stability and structure of the flame. Kadowaki et al. [3] performed 2-D numerical simulations of a TD unstable hydrogen-air flame and also found that the size of the computational significantly influences the flame dynamics. Berger et al. [4] increased the domain size even further in attempt to rule out these confinement effects on the flame front dynamics. They found that for a sufficiently large domain size, the flame consumption speed becomes independent of the domain size. In later work, Berger et al. [5] quantified the contribution of each instability mechanism (DL and TD) which confirms the destabilizing effect of the TD instabilities for lean hydrogen-air mixtures, in line with the theory of Matalon et al. [10]. Creta et al. [6] showed that hydrodynamic (DL) instabilities lead to large-scale cusp-like flame structures, while TD instabilities introduce small-scale corrugations. The interaction of these instabilities results in complex flame morphologies and enhanced propagation speeds. In later work of Kadowaki et al. [7] they performed 2- and 3-D numerical simulations of TD unstable cellular hydrogen-air flames and found that in 3-D the increment of flame surface area and flame velocity is about twice that of 2-D cellular flames, which suggests that the TD effects are even more prominent in 3-D. This has also recently been confirmed by Wen et al. [8], who did a large-scale 3-D numerical simulation of a laminar lean premixed hydrogen-air flame. They quantified the differences between 2- and 3-D TD unstable flames for sufficiently large domain sizes to rule out the influence of domain size, and found that for 3-D flames the reactivity is enhanced due to higher peak curvature values. Howarth et al. [11, 12] performed 2- and 3-D simulations of lean TD unstable hydrogen flames and identified a single empirical instability parameter $\omega_{2}$ that relates the multi-dimensional unstable flame speed and thickness to their laminar one-dimensional values.
@@ 2.2 FGM Control Variables
Since the Lewis number of hydrogen is significantly lower than unity, preferential diffusion effects result in local changes in mixture composition. To address this, a second control variable $\phi_{2}$ is introduced: the elementary mixture fraction of hydrogen,
@@
with $c_{Z,i}$ being the mass fraction of hydrogen elements in species $i$. In most engineering applications, heat losses play an important role. Therefore, a third control variable $\phi_{3}$ is introduced: the enthalpy,
@@ 2.3 Preferential Diffusion Analysis
The preferential diffusion term in Eq. (9) can be computed by storing the diffusion coefficients and mass fractions of all chemical species present in the chemical mechanism in the lookup table. During the CFD calculation, these coefficients and mass fractions are retrieved using the control variables, and the diffusion fluxes of all species are then calculated to construct the preferential diffusion term. However, for large chemical mechanisms this results in a significant increase in memory occupation of the lookup table and extra computational overhead due to retrieving all variables and computing the species gradients. A simpler method could be to include solely the major species which have a significant contribution to the total preferential diffusion flux of the control variable, which are species that either have $\text{Le}\neq 1$ or a large thermal diffusion coefficient $D_{i}^{T}$. To analyse this, the contributions of each chemical species to the preferential diffusion flux of the mixture fraction and enthalpy, i.e., the terms within the summation of the right term in Eq. (9), are plotted in Figure 1. The results are obtained via a one-dimensional flame simulation of a hydrogen-air mixture at an equivalence ratio of $\phi=0.4$ and at ambient pressure ($101325$ Pa) and temperature ($300$ K). The chemical mechanism used is by Burke et al. [32] and the simulations are performed using an in-house one-dimensional flamelet code Chem1D [33].
@@
Figure 1: The contributions of each chemical species to the preferential diffusion flux of mixture fraction (a) and enthalpy (b), and the temperature profile obtained from a one-dimensional premixed hydrogen-air flame simulation at an equivalence ratio of $\phi=0.4$.
@@
In both plots of Figure 1, the dotted line indicates the point where the temperature is equal to 400 K, defined as the zero-position. In Figure 1(a), it can be observed that there is a positive flux of mixture fraction due to the diffusion of hydrogen before zero-position, which is a result of hydrogen’s low Lewis number. From Figure 1(a) it can be concluded that the preferential diffusion flux of mixture fraction is mainly governed by H2 ($\approx 94\%$), less by H2O ($\approx 4\%$) and slightly by H ($\approx 1\%$). All other species seem to have no significant influence on the preferential diffusion of mixture fraction. The species contributions to the preferential diffusion of enthalpy, plotted in Figure 1(b), seem to be more diverse. The species H, H2 and H2O seem to be the most significantly contributing species ($\approx 90\%$ combined) with all three having a similar magnitude, while species such as O2, O and OH having a slight contribution ($\approx 7\%$) as well. The progress variable is excluded from this analysis as it is constructed solely from H2 and H2O, which results in $\mathbf{J}_{\mathcal{Y},i}=0$ except for H2 and H2O.
@@ 2.4 Manifolds & Lookup
In this way, we attempt to approximate the three-dimensional divergence of the diffusion flux of the remaining species using the one-dimensional divergence in the flamelets.
@@
Note that the small contribution of the H-radical to the mixture fraction preferential diffusion, which can be seen in Figure 1(a), is now also included directly in the preferential diffusion term. The small contributions of the remaining species to the preferential diffusion of enthalpy, mainly O2, O and OH as seen in Figure 1(b), are gathered in the closure source term $S_{h}$. Note that for the progress variable, the contribution of the H-radical is zero, since the progress variable is constructed from H2 and H2O only. The hypothesis is that by including the contribution of the H-radical directly in the preferential diffusion of enthalpy, cross-diffusion and the TD instabilities due to preferential diffusion and curvature can be modeled with significantly more accuracy using FGM B2 compared to FGM B1.
@@ 3.1 Setup
The simulation setup is similar to the work of Berger et al. [4]. The two-dimensional domain has a height $L_{y}=200l_{F}$, where $l_{F}$ represents the laminar flame thickness. In the work of Berger et al. [4], the domain width $L_{x}$ is varied and it is shown that the average flame consumption speed stabilizes for $\log_{10}(L_{x}/l_{F})>2$, which corresponds to a domain width of $L_{x}>100l_{F}$. Therefore, a domain width of $100l_{F}$ is chosen. The initial flame front is perturbed with a sinusoidal disturbance with a wavelength $\lambda=10l_{F}$. The considered fuel mixture is hydrogen in air at an equivalence ratio $\phi=0.4$ at an ambient pressure of $101325$ Pa and with a temperature of $300$ K. At these conditions, $l_{F}$ has a value of approximately 0.67 mm. A uniform inlet velocity is chosen to balance the flame propagation speed and keep the flame front in the computational domain. The velocity has a value of $0.70$ m/s and is kept constant for all simulated cases. For both the detailed chemistry calculation and the flamelets for FGM table generation, the chemical mechanism by Burke et al. [32] is used. All simulations run with mixture averaged diffusion coefficients and thermal (Soret) diffusion. The total simulation time for all cases is approximately 150 $\tau_{F}$, where $\tau_{F}$ is the laminar flame time, defined as $\tau_{F}=l_{F}/s_{L}$. Here, $s_{L}$ denotes the laminar flame speed, which has a value of 204.85 mm/s at these conditions, resulting in $\tau_{F}=3.26$ ms. Note that the inlet velocity is more than 3 times higher than the laminar burning velocity of this mixture. This is because the increased flame surface area due to the unstable wrinkled flame front results in a much larger burning rate.
@@ 3.2 A-priori Validation of the Manifolds
Figure 2: Contour plot of the normalized temperature $\Theta=(T-T_{ad})/(T_{u}-T_{ad})$ for the detailed chemistry (DC) reference case.
@@
Figure 3: A-priori analysis of manifold A and manifold B using the DC data. The left column show DC values normalized by their 1D maximum values. The right two columns show the relative error of the manifolds compared to the DC values, defined as $\epsilon_{\psi}=(\psi_{\text{DC}}-\psi_{\text{FGM}})/\text{max}(\psi_{\text{DC}})\cdot 100\%$.
In Figure 3 the a-priori analysis of the manifolds is visualized. The normalized progress variable source term and the normalized H-radical mass fraction obtained from the DC reference calculation are plotted in the upper left and lower left figures, respectively. The control variables associated with each manifold are obtained from the DC fields and a lookup of the variables in consideration is performed. The difference between the manifold lookup variables and the DC values is quantified by the relative error $\epsilon_{\psi}$. The magnitude of this error is integrated over area enclosed by the dashed lines, which defined as where the normalized progress variable source term is higher than a threshold value of $\frac{\dot{\omega}_{\mathcal{Y}}}{\text{max}(\dot{\omega}_{\mathcal{Y},\text{1D}})}=0.165$. This integrated error value is added to the plot as $|\bar{\epsilon}|_{\psi}$. It can be seen that by adding enthalpy as an extra dimension to the manifold, both variables are represented with slightly higher accuracy. The contour plot of the progress variable source term error $\epsilon_{\dot{\omega}_{\mathcal{Y}}}$ does not seem to show a significant improvement, but the integrated error magnitude $|\bar{\epsilon}|_{\dot{\omega}_{\mathcal{Y}}}$ does show a slight improvement of manifold B with respect to manifold A. The surface plots of the H-radical error $\epsilon_{Y_{H}}$ seem to show a slightly more significant error reduction in the flame tip region, which is also reflected by a reduced integrated error magnitude $|\bar{\epsilon}|_{\dot{\omega}_{\mathcal{Y}}}$. Nonetheless, it can be concluded that both manifolds show no significant deviations from the DC reference case when compared in an a-priori sense.
@@ 3.3 Linear Stability Analysis
Figure 4: A comparison of the non-dimensional dispersion relation between detailed chemistry (DC) and the FGM models. The dots represent the calculated dimensionless growth rates and the lines represent a third order polynomial fit through these points.
The exponential growth of small perturbations ultimately leads to the typical cell structures found in TD unstable flames. Therefore, the growth of perturbations for different wave lengths is studied by performing a linear stability analysis (LSA), which has proven to be a useful tool in investigating flame front instabilities [4, 5]. The planar flame front is disturbed with a small initial perturbation of $A_{0}=0.04l_{F}$ for varying wave lengths $\lambda$, similar as in the work of Berger et al. [4] and Böttler et al. [28]. The lateral domain size is adjusted such that exactly one wave length fits in the domain. For each wave length, the growth rate $\omega=d(\ln{A(t)}/A_{0})/dt$ is calculated. Figure 4 shows the dimensionless growth rate for varying dimensionless wave numbers. A positive number indicates an unstable exponential growth of the perturbation, while a negative number indicates an exponential dampening of the perturbation.
[…]
@@ 3.4 A-posteriori Unstable Flame Front Simulation
Figure 5: Contour plot of the normalized temperature $\Theta=(T-T_{\text{ad}})/(T_{\text{u}}-T_{\text{ad}})$ for the compared models.
@@
In Figure 5, the DC calculation is compared with the FGM models by lining up snapshots of the normalized temperature contours at $t\approx 70\tau_{F}$. The temperature is normalized using unburnt temperature $T_{\text{u}}$ of 300 K and the adiabatic flame temperature $T_{\text{ad}}$ of the inlet mixture ($\phi=4.0$). It can be observed that in post-flame regions where the flame is positively curved, i.e., convex towards the incoming flow, the normalized temperature increases to values above unity. Due to preferential diffusion, relatively more H2 diffuses toward these regions, leading to a richer mixture and consequently a super-adiabatic temperature. In contrast, in negatively curved regions, i.e., concave towards the incoming flow, the mixture becomes leaner, leading to a trail of sub-adiabatic temperatures. This effect is visible in all models. While investigating the shape of the wrinkled flame front, DC, FGM model A and B2 appear to show the characteristic flame tips, similar to those reported in several literature sources [1, 2, 3, 4, 5, 6, 7, 8]. These flame tips consist of a single flame cusp enclosed by two flame bulges. This structure remains stable and tilts towards an incline with respect to the incoming flow. This results in a lateral displacement in positive or negative $x$-direction depending on the flame tip orientation, until the flame tip reaches a different part of the flame front and merges. FGM model B1 deviates from this dynamic behavior. The typical flame tip structure, a stable single cusp enclosed with two bulges characteristic for the other models, can not be observed in the results of FGM B1. Instead, unstable, single flame fingers grow fast and move rapidly downwards. Additionally, there seems to be no lateral movement of the flame tips. Furthermore, the temperature seems to be higher in these large flame fingers, which could be attributed to the high positive curvature in the tip.
@@
To further highlight the differences in dynamic behavior between the models, snapshots of the flame fronts at several time instances are plotted in Figure 6. The black lines represent the flame fronts, defined as the $\Theta=0.5$ isocontour. The flame front that enters the figure at $y^{\prime}/l_{F}=0$ corresponds to $t=38\tau_{F}$. For each time instance the flame front is displaced downwards with a value $\Delta y^{\prime}=U\Delta t$, where $U$ represents the inflow velocity and $\Delta t\approx 0.613\tau_{F}$. This is done to facilitate a clear distinction between the flame fronts. Some of the flame cusps for the models DC, FGM A, and FGM B2, are tracked by the dashed blue and red lines which represent the flame cusps moving in positive and negative x-direction, respectively. The colored dots and crosses represent the locations of emergence and disappearance of the flame cusps, respectively. In the figure for FGM B1, red dots mark the points where some flame fingers emerge, and their movement is shown with red arrows.
@@
Figure 6: Flame front visualisation to analyse the dynamic behaviour of the different models. The solid black lines indicate the flame front at different time instances, defined as the $\Theta=0.5$ normalized temperature isocontour. The colored dashed lines show the movement of the flame cusps for models DC, FGM A, and FGM B2, while the red arrows indicate the movement of a single flame finger for FGM B1.
@@
For DC, FGM A and FGM B2, the flame tips that enter at $y^{\prime}/l_{F}=0$ are either already inclined, or proceed to tilt towards an angle with respect to the incoming flow. New flame cusps seem to emerge from the existing laterally displacing flame tips, where two bulges on each side expand, resulting in a new flame tip structure. The flame tips get destroyed by collision with other parts of the flame front due to their lateral displacement. This behavior is in line with the findings of Berger et al. [4]. In contrast to the other models, FGM B1 does not show this behavior. The flame finger formation begins as expected, with a single cusp forming at the red dot and two enclosing bulges starting to grow. However, instead of developing into a stable structure, one bulge grows faster than the other, leading to the disappearance of the smaller bulge. The larger bulge continues to grow, following the red line, but grows too large to remain stable. A new cusp forms, and the process repeats. Due to the rapid growth and downward motion of the flame fingers, the entire flame front moves downward much faster compared to the other models.
@@
This is also evident when looking at the total progress variable chemical production rate over the simulation time in Figure 7. The chemical production rate is computed by integrating the progress variable source term over the domain. The thin lines represent the instantaneous values and the thick lines are the running time-averaged values. One can see that the time averaged progress variable production rates of the DC, FGM A and FGM B2 are stabilizing towards a similar value, while FGM B1 immediately increases to a significantly higher value. Because the inflow velocity is kept constant for all simulations, this results in the flame propagating much faster towards the inlet for FGM B1. Therefore, the line corresponding to FGM B1 is cut-off at $t\approx 87\tau_{F}$, as the flame front reaches the inflow boundary at $y/l_{F}=0$ and stabilizes there. At this location the flame front dynamics are influenced by the inflow boundary and therefore the rest of the FGM B1 simulation is discarded.

### v2

@@ 1 Introduction
Lean laminar TD unstable hydrogen-air flames have been investigated by numerous sources in the literature [1, 2, 3, 4, 5, 6, 7, 8]. Altantzis et al. [1, 2] studied the initial linear growth of superimposed perturbations and the long-term nonlinear evolution of instabilities by 2-D numerical simulations with single-step chemistry and detailed transport for varying domain sizes. They show that positively-curved segments of the flame experience increased reactivity, while the reactivity in negatively-curved segments decreases. Furthermore, that the domain size significantly influences the stability and structure of the flame. Kadowaki et al. [3] performed 2-D numerical simulations of a TD unstable hydrogen-air flame and also found that the size of the computational significantly influences the flame dynamics. Berger et al. [4] increased the domain size even further in attempt to rule out these confinement effects on the flame front dynamics. They found that for a sufficiently large domain size, the flame consumption speed becomes independent of the domain size. In later work, Berger et al. [5] quantified the contribution of each instability mechanism (DL and TD) which confirms the destabilizing effect of the TD instabilities for lean hydrogen-air mixtures, in line with the theory of Matalon et al. [10]. Creta et al. [6] showed that hydrodynamic (DL) instabilities lead to large-scale cusp-like flame structures, while TD instabilities introduce small-scale corrugations. The interaction of these instabilities results in complex flame morphologies and enhanced propagation speeds. In later work of Kadowaki et al. [7] they performed 2- and 3-D numerical simulations of TD unstable cellular hydrogen-air flames and found that in 3-D the increment of flame surface area and flame velocity is about twice that of 2-D cellular flames, which suggests that the TD effects are even more prominent in 3-D. This has also recently been confirmed by Wen et al. [8], who did a large-scale 3-D numerical simulation of a laminar lean premixed hydrogen-air flame. They quantified the differences between 2- and 3-D TD unstable flames for sufficiently large domain sizes to rule out the influence of domain size, and found that for 3-D flames the reactivity is enhanced due to higher peak curvature values. Howarth et al. [11, 12] performed 2- and 3-D simulations of lean TD unstable hydrogen flames and identified a single empirical instability parameter $\omega_{2}$ that relates the multi-dimensional unstable flame speed and thickness to their laminar one-dimensional values. Aspden et al. [13, 14] investigated diffusive effects in turbulent lean premixed methane and hydrogen flames, and found that the H-radical plays a particularly prominent role in hydrogen flames. The rapid diffusion of atomic hydrogen has a strong impact on flame structure; their three-dimensional DNS results show that neglecting this effect leads to thinner flames, higher peak reaction rates, and a significantly altered coupling between flame curvature and local reactivity.
@@ 2.2 FGM Control Variables
Since the Lewis number of hydrogen is significantly lower than unity, preferential diffusion effects result in local changes in mixture composition. To address this, a second control variable $\phi_{2}$ is introduced: the element based mixture fraction of the hydrogen element,
@@
with $c_{Z,i}$ being the mass fraction of the hydrogen element in species $i$. In most engineering applications, heat losses play an important role. Therefore, a third control variable $\phi_{3}$ is introduced: the enthalpy,
@@ 2.3 Preferential Diffusion Analysis
The preferential diffusion term in Eq. (9) can be computed by storing the diffusion coefficients and mass fractions of all chemical species present in the chemical mechanism in the lookup table. During the CFD calculation, these coefficients and mass fractions are retrieved using the control variables, and the diffusion fluxes of all species are then calculated to construct the preferential diffusion term. However, for large chemical mechanisms this results in a significant increase in memory occupation of the lookup table and extra computational overhead due to retrieving all variables and computing the species gradients. A simpler method could be to include solely the major species which have a significant contribution to the total preferential diffusion flux of the control variable, which are species that either have $\text{Le}\neq 1$ or a large thermal diffusion coefficient $D_{i}^{T}$. To analyse this, the contributions of each chemical species to the preferential diffusion flux of the mixture fraction and enthalpy, i.e., the terms within the summation of the right term in Eq. (9), are plotted in Figure 1. The results are obtained from one-dimensional flame simulations at ambient pressure ($101325$ Pa) and temperature ($300$ K). The mixture consists of hydrogen in air at two equivalence ratios: $\phi=0.4$ and $0.7$. These cases are included to illustrate that the relative flux contributions of individual species remain similar across the range of equivalence ratios that can occur due to thermo-diffusive instabilities. The chemical mechanism used is by Burke et al. [34] and the simulations are performed using an in-house one-dimensional flamelet code Chem1D [35].
(a) $\phi=0.4$
(b) $\phi=0.4$
(c) $\phi=0.7$
(d) $\phi=0.7$
@@
Figure 1: The contributions of each chemical species to the preferential diffusion flux of mixture fraction (a,c) and enthalpy (b,d), and the temperature profile obtained from a one-dimensional premixed hydrogen-air flame simulation at an equivalence ratio of $\phi=0.4$ (a,b) and $\phi=0.7$ (c,d).
@@
In all plots of Figure 1, the dotted vertical line indicates the point where the temperature is equal to 400 K, defined as the zero-position. In Figure 1(a) and 1(c), it can be observed that there is a positive flux of mixture fraction due to the diffusion of hydrogen before zero-position, which is a result of hydrogen’s low Lewis number. From Figures 1(a) and 1(c), it can be concluded that the preferential diffusion flux of the mixture fraction is predominantly governed by H2, contributing approximately 94% at $\phi=0.4$ and 93% at $\phi=0.7$. The contributions of H2O and H are significantly smaller: around 4% and 1%, respectively, at $\phi=0.4$, and 2% and 5% at $\phi=0.4$. All other species seem to have no significant influence on the preferential diffusion of mixture fraction. The species contributions to the preferential diffusion of enthalpy, plotted in Figure 1(b) and 1(d), seem to be more diverse. The species H, H2 and H2O seem to be the most significantly contributing species ($\approx 90\%$ for both $\phi$), while species such as O2, O and OH having a slight contribution ($\approx 7\%$) as well. The progress variable is excluded from this analysis as it is constructed solely from H2 and H2O, which results in $\mathbf{J}_{\mathcal{Y},i}=0$ except for H2 and H2O.
@@ 2.4 Manifolds & Lookup
In this way, we attempt to approximate the three-dimensional divergence of the diffusion flux of the remaining species using the one-dimensional divergence in the flamelets. This introduces an error, but following Figure 1, it will be assumed to be negligible.
@@
Note that the small contribution of the H-radical to the mixture fraction preferential diffusion, which can be seen in Figure 1(a) and 1(c), is now also included directly in the preferential diffusion term. The small contributions of the remaining species to the preferential diffusion of enthalpy, mainly O2, O and OH as seen in Figure 1(b) and 1(d), are gathered in the closure source term $S_{h}$. Note that for the progress variable, the contribution of the H-radical is zero, since the progress variable is constructed from H2 and H2O only. The hypothesis is that by including the contribution of the H-radical directly in the preferential diffusion of enthalpy, cross-diffusion and the TD instabilities due to preferential diffusion and curvature can be modeled with significantly more accuracy using FGM B2 compared to FGM B1.
@@ 3.1 Setup
The simulation setup is similar to the work of Berger et al. [4]. The two-dimensional domain has a height $L_{y}=200l_{F}$, where $l_{F}$ represents the unstretched laminar flame thickness defined as $l_{F}=\left(T_{b}-T_{u}\right)/\max\left(\frac{dT}{dx}\right)$. In the work of Berger et al. [4], the domain width $L_{x}$ is varied and it is shown that the average flame consumption speed stabilizes for $\log_{10}(L_{x}/l_{F})>2$, which corresponds to a domain width of $L_{x}>100l_{F}$. Therefore, a domain width of $100l_{F}$ is chosen. The mesh resolution was set to 20 cells in the unstretched laminar flame front, i.e. $\Delta x=l_{F}/20$. The authors found that further refinement beyond this resolution did not affect the dynamic behavior for any of the models considered. The initial flame front is perturbed with a sinusoidal disturbance with a wavelength $\lambda=10l_{F}$. The considered fuel mixture is hydrogen in air at an equivalence ratio $\phi=0.4$ at an ambient pressure of $101325$ Pa and with a temperature of $300$ K. At these conditions, $l_{F}$ has a value of approximately 0.67 mm. A uniform inlet velocity is chosen to balance the flame propagation speed and keep the flame front in the computational domain. The velocity has a value of $0.70$ m/s and is kept constant for all simulated cases. For both the detailed chemistry calculation and the flamelets for FGM table generation, the chemical mechanism by Burke et al. [34] is used. All simulations run with mixture averaged diffusion coefficients and thermal (Soret) diffusion. The total simulation time for all cases is approximately 150 $\tau_{F}$, where $\tau_{F}$ is the laminar flame time, defined as $\tau_{F}=l_{F}/s_{L}$. Here, $s_{L}$ denotes the unstretched laminar flame speed, which has a value of 204.85 mm/s at these conditions, resulting in $\tau_{F}=3.26$ ms. Note that the inlet velocity is more than 3 times higher than the laminar burning velocity of this mixture. This is because the increased flame surface area due to the unstable wrinkled flame front results in a much larger burning rate. The FGM model’s computational cost is about an order of magnitude higher than that of detailed chemistry (DC), while the difference between individual FGM models is significantly smaller.
@@ 3.2 A-priori Validation of the Manifolds
Figure 2: Contour plot of the normalized temperature $\Theta=(T-T_{\mathrm{u}})/(T_{\mathrm{ad}}-T_{\mathrm{u}})$ for the detailed chemistry (DC) reference case.
@@
Figure 3: A-priori analysis of manifold A and manifold B using the DC data. The left column show DC values normalized by their 1D maximum values. The right two columns show the normalized error of the manifolds compared to the DC values, defined as $\epsilon_{\psi}=(\psi_{\text{DC}}-\psi_{\text{FGM}})/\text{max}(\psi_{\text{DC}})\cdot 100\%$.
In Figure 3 the a-priori analysis of the manifolds is visualized. The normalized progress variable source term and the normalized H-radical mass fraction obtained from the DC reference calculation are plotted in the upper left and lower left figures, respectively. The control variables associated with each manifold are obtained from the DC fields and a lookup of the variables in consideration is performed. The difference between the manifold lookup variables and the DC values is quantified by the normalized error $\epsilon_{\psi}$. The magnitude of this error is integrated over and divided by the area enclosed by the dashed lines, which defined as where the normalized progress variable source term is higher than a threshold value of $\frac{\dot{\omega}_{\mathcal{Y}}}{\text{max}(\dot{\omega}_{\mathcal{Y},\text{1D}})}=0.165$. This average error value is added to the plot as $|\bar{\epsilon}|_{\psi}$. It can be seen that by adding enthalpy as an extra dimension to the manifold, both variables are represented with slightly higher accuracy. The contour plot of the progress variable source term error $\epsilon_{\dot{\omega}_{\mathcal{Y}}}$ does not seem to show a significant improvement, but the average error magnitude $|\bar{\epsilon}|_{\dot{\omega}_{\mathcal{Y}}}$ does show a slight improvement of manifold B with respect to manifold A. The surface plots of the H-radical error $\epsilon_{Y_{H}}$ seem to show a slightly more significant error reduction in the flame tip region, which is also reflected by a reduced average error magnitude $|\bar{\epsilon}|_{\dot{\omega}_{\mathcal{Y}}}$. Nonetheless, it can be concluded that both manifolds show no significant deviations from the DC reference case when compared in an a-priori sense.
@@ 3.3 Linear Stability Analysis
Figure 4: A comparison of the non-dimensional dispersion relation between detailed chemistry (DC) and the FGM models. The dots represent the calculated dimensionless growth rates and the lines represent a third order polynomial fit through these points. The black dash-dotted line represents the Darrieus-Landau growth rate obtained from the work of Matalon [36].
The exponential growth of small perturbations ultimately leads to the typical cell structures found in TD unstable flames. Therefore, the growth of perturbations for different wave lengths is studied by performing a linear stability analysis (LSA), which has proven to be a useful tool in investigating flame front instabilities [4, 5]. The planar flame front is disturbed with a small initial perturbation of $A_{0}=0.04l_{F}$ for varying wave lengths $\lambda$, similar as in the work of Berger et al. [4] and Böttler et al. [30]. The lateral domain size is adjusted such that exactly one wave length fits in the domain, and to ensure accurate results, a high mesh resolution of 100 cells in the laminar flame front is used, i.e. $\Delta x=l_{F}/100$. For each wave length, the growth rate $\omega=d(\ln{A(t)}/A_{0})/dt$ is calculated. Figure 4 shows the dimensionless growth rate for varying dimensionless wave numbers. A positive number indicates an unstable exponential growth of the perturbation, while a negative number indicates an exponential dampening of the perturbation.
The global behavior seems to be consistent, showing a growth of perturbations when the wavelengths are not too small, and a peak around $4l_{F}$. Additionally, all FGM models show unstable growth of perturbations at smaller wavelengths compared to the DC reference case. However, FGM model B1 seems to deviate the most from DC, showing similar results to the FPV-$h$ model presented in the work of Böttler et al. [30]. This can be due to the fact that these models are in their core very similar. The FPV-$h$ model does not explicitly take into account the contribution of the H-radical to the preferential diffusion of enthalpy, since their major species selection only consists of H2, O2 and H2O. By explicitly including the H-radical in the enthalpy preferential diffusion term, as done for FGM model B2, a significant improvement can be observed. FGM B2 also shows a slight improvement compared to the FPV-$\kappa_{c}$ model proposed by Böttler et al. [30], which attempts to include curvature effects in the manifold. FGM A also appears to perform well, which is expected given the small errors observed in the a-priori analysis in Section 3.2. This indicates that the main deviations arise not from the manifolds themselves but from the solution of the control variable transport equations. Specifically, our analysis suggests that discrepancies in the predicted control variables, mainly influenced by the preferential diffusion formulation as well as the general FGM assumption, play a significant role in altering the flame’s stability characteristics. To further analyze the models, an unstable flame front is simulated for a longer time to study the subsequent non-linear behavior.
@@ 3.4 A-posteriori Unstable Flame Front Simulation
Figure 5: Contour plot of the normalized temperature $\Theta=(T-T_{\mathrm{u}})/(T_{\mathrm{ad}}-T_{\mathrm{u}})$ for the compared models.
@@
In Figure 5, the DC calculation is compared with the FGM models by lining up snapshots of the normalized temperature contours at $t\approx 67.4\tau_{F}$. The temperature is normalized using unburnt temperature $T_{\text{u}}$ of 300 K and the adiabatic flame temperature $T_{\text{ad}}$ of the inlet mixture ($\phi=4.0$). It can be observed that in post-flame regions where the flame is positively curved, i.e., convex towards the incoming flow, the normalized temperature increases to values above unity. Due to preferential diffusion, relatively more H2 diffuses toward these regions, leading to a richer mixture and consequently a super-adiabatic temperature. In contrast, in negatively curved regions, i.e., concave towards the incoming flow, the mixture becomes leaner, leading to a trail of sub-adiabatic temperatures. This effect is visible in all models. While investigating the shape of the wrinkled flame front, DC, FGM model A and B2 appear to show the characteristic flame tips, similar to those reported in several literature sources [1, 2, 3, 4, 5, 6, 7, 8]. These flame tips consist of a single flame cusp enclosed by two flame bulges. This structure remains stable and tilts towards an incline with respect to the incoming flow. This results in a lateral displacement in positive or negative $x$-direction depending on the flame tip orientation, until the flame tip reaches a different part of the flame front and merges. FGM model B1 deviates from this dynamic behavior. The characteristic flame tip structure is still present, but it does not exhibit the dynamic instability or lateral tilting observed in the other models. Instead, a single structure remains stable, oriented perpendicular to the flow, and propagates upstream, while smaller secondary flame tips are suppressed and rarely emerge.
@@
To further highlight the differences in dynamic behavior between the models, snapshots of the flame fronts at several time instances are plotted in Figure 6. The black lines represent the flame fronts, defined as the $\Theta=0.5$ isocontour. The flame front that enters the figure at $y^{\prime}/l_{F}=0$ corresponds to $t=38\tau_{F}$. For each time instance the flame front is displaced downwards with a value $\Delta y^{\prime}=U\Delta t$, where $U$ represents the inflow velocity and $\Delta t\approx 0.613\tau_{F}$. This is done to facilitate a clear distinction between the flame fronts. Some of the flame cusps for all the models, are tracked by the dashed blue and red lines which represent the flame cusps moving in positive and negative x-direction, respectively. The colored dots and crosses represent the locations of emergence and disappearance of the flame cusps, respectively.
@@
Figure 6: Flame front visualisation to analyse the dynamic behaviour of the different models. The solid black lines indicate the flame front at different time instances, defined as the $\Theta=0.5$ normalized temperature isocontour. The colored dashed lines show the movement of the flame cusps for all the models.
@@
For DC, FGM A and FGM B2, the flame tips that enter at $y^{\prime}/l_{F}=0$ are either already inclined, or proceed to tilt towards an angle with respect to the incoming flow. New flame cusps seem to emerge from the existing laterally displacing flame tips, where two bulges on each side expand, resulting in a new flame tip structure. The flame tips get destroyed by collision with other parts of the flame front due to their lateral displacement. This behavior is in line with the findings of Berger et al. [4]. In contrast to the other models, FGM B1 does not show this behavior. A single flame finger enters the domain and continues to propagate downstream while maintaining a stable orientation perpendicular to the incoming flow. Although smaller secondary flame tips occasionally emerge from the sides of this dominant structure, they are destroyed upon interaction with the main flame tip. Due to this persistent dominant flame tip, the flame front moves downwards significantly faster compared to the other models.
@@
This is also evident when looking at the total progress variable chemical production rate over the simulation time in Figure 7. The chemical production rate is computed by integrating the progress variable source term over the domain. The thin lines represent the instantaneous values and the thick lines the averaged values over the window $t/\tau_{F}\in[40,85]$. An averaging start time of $t/\tau_{F}=40$ is selected to ensure that initialization effects do not affect the averaged values.
One can see that the averaged progress variable production rates in Figure 7 of the DC, FGM A and FGM B2 are significantly lower than of FGM B1. Because the inflow velocity is kept constant for all simulations, this results in the flame propagating much faster towards the inlet for FGM B1. Notably, model FGM A also exhibits a slight overprediction in burning rate, suggesting that the enthalpy itself may contribute to an accurate prediction of thermo-diffusive instabilities.